OpenStudy (lena772):

Which of the following best defines 4^(2/3)? Square root of 16 Cube root of 4 Square root of 4 Cube root of 16

OpenStudy (lena772):

one of these again I can't even

OpenStudy (kc_kennylau):

Try calculating the values of the five expressions :D

OpenStudy (agent0smith):

\[\Large 4^{2/3} = 4^{2*1/3} = (4^2)^{1/3}\]

OpenStudy (lena772):

Is it cubed root of 4?

OpenStudy (kc_kennylau):

no :/ Try again :D

OpenStudy (lena772):

shi...taki

OpenStudy (agent0smith):

mushrooms?

OpenStudy (lena772):

cube root of 16?

OpenStudy (kc_kennylau):

yep :D

OpenStudy (lena772):

-does hula dance-

OpenStudy (kc_kennylau):

lol :D

OpenStudy (agent0smith):

:D